# LyricWhiz: Robust Multilingual Zero-shot Lyrics Transcription by Whispering to ChatGPT

### v1

@@ Abstract
We introduce LyricWhiz, a robust, multilingual, and zero-shot automatic lyrics transcription method achieving state-of-the-art performance on various lyrics transcription datasets, even in challenging genres such as rock and metal. Our novel, training-free approach utilizes Whisper, a weakly supervised robust speech recognition model, and GPT-4, today’s most performant chat-based large language model. In the proposed method, Whisper functions as the “ear” by transcribing the audio, while GPT-4 serves as the “brain,” acting as an annotator with a strong performance for contextualized output selection and correction. Our experiments show that LyricWhiz significantly reduces Word Error Rate compared to existing methods in English and can effectively transcribe lyrics across multiple languages. Furthermore, we use LyricWhiz to create the first publicly available, large-scale, multilingual lyrics transcription dataset with a CC-BY-NC-SA copyright license, based on MTG-Jamendo, and offer a human-annotated subset for noise level estimation and evaluation. We anticipate that our proposed method and dataset will advance the development of multilingual lyrics transcription, a challenging and emerging task.
@@ 1 Introduction
In this paper, we present LyricWhiz, a novel method for automatic lyrics transcription. LyricWhiz surpasses existing methods on various ALT datasets, resulting in a significant reduction in WER for English lyrics and providing accurate transcription results across multiple languages. Our system is robust, multilingual, and training-free. To achieve these results, we combined two powerful models from their respective domains as shown in Figure 1: Whisper, a weakly supervised speech transcription model, and GPT-4, a large language model (LLM) from the ChatGPT family. Whisper acts as the “ear” while GPT-4 serves as the “brain” by providing contextualized output selection and correction with strong performance [9]. We further use LyricWhiz to build a multilingual lyrics dataset, named MulJam, which to our knowledge, is the first large-scale, multilingual lyrics transcription dataset without copyright-related issues.
@@ 2.1 Automatic Lyrics Transcription
To date, there are no datasets for multilingual ALT, which makes transfer learning across languages, as commonly done for ASR, infeasible for ALT. We investigate the large-scale weakly supervised ASR model and discover that we can directly apply it to multilingual ALT for zero-shot prediction and also achieve impressive results. This further motivates us to leverage large-scale open datasets in other MIR tasks to build the first multilingual ALT dataset to unlock the potential of multilingual ALT models.
@@ 2.3 Chat-based Large Language Models
c) Drawing inspiration from the recent success of employing communicative agents [42] and generative agents [43] to generate new instruction corpora and activate CoT prompting, we assign ChatGPT the role of transcription post-processor, allowing ChatGPT to access more domain-specific knowledge and capabilities. The importance of a well-designed CoT [44], which effectively divides a comprehensive task into several phases and designs specific prompts for each phase, is widely acknowledged for enhancing LLM performance. LyricWhiz also proposes a concise CoT strategy, depicted in Section 3.2.
@@ 3.1 Whisper as Zero-shot Lyrics Transcriptor
Task: As a GPT-4 based lyrics transcription post-processor, your task is to analyze multiple ASR model-generated versions of a song’s lyrics and determine the most accurate version closest to the true lyrics. Also filter out invalid lyrics when all predictions are nonsense.
Input: The input is in JSON format:
{“prediction_1”: “line1;line2;…”, …}
Output: Your output must be strictly in readable JSON format without any extra text:
“reasons”: “reason1;reason2;…”,
“closest_prediction”: <key_of_prediction>
“output”: “line1;line2…”
Requirements: For the "reasons" field, you have to provide a reason for the choice of the "closest_prediction" field. For the "closest_prediction" field, choose the prediction key that is closest to the true lyrics. Only when all predictions greatly differ from each other or are completely nonsense or meaningless, which means that none of the predictions is valid, fill in "None" in this field. For the "output" field, you need to output the final lyrics of closest_prediction. If the "closest_prediction" field is "None", you should also output "None" in this field. The language of the input lyrics is English.
@@ 3.3 Multilingual Lyrics Transcription Dataset
Following the construction process outlined above, we ultimately obtained a multilingual lyric transcription dataset, MulJam, consisting of 6,031 songs with 182,429 lines and a total duration of 381.9 hours. The dataset’s statistical information and comparisons with existing ALT datasets are presented in Table 2. To our best knowledge, MulJam is the first publicly available large-scale dataset for multilingual lyrics transcription without copyright restrictions. While DALI [7] is another large-scale music dataset featuring multilingual lyrics, its restricted access and strict licensing requirements limit its applicability for downstream tasks. In contrast, MulJam is free from copyright-related constraints and can be utilized without approval, as the audio can be legally downloaded directly from public sources without the need for approval, making it easily accessible. This even includes audio that is permitted for use in the development of commercial software. Researchers are permitted to legally modify our dataset for derivative works and redistribution, provided they cite our work and adhere to the CC BY-NC-SA license. Furthermore, in contrast to the imbalanced language distribution in DALI, where English songs account for over $80\%$ of the total songs, our dataset includes a greater proportion of songs in other languages, which is advantageous for multilingual lyrics transcription.

### v4

@@ Abstract
We introduce LyricWhiz, a robust, multilingual, and zero-shot automatic lyrics transcription method achieving state-of-the-art performance on various lyrics transcription datasets, even in challenging genres such as rock and metal. Our novel, training-free approach utilizes Whisper, a weakly supervised robust speech recognition model, and GPT-4, today’s most performant chat-based large language model. In the proposed method, Whisper functions as the “ear” by transcribing the audio, while GPT-4 serves as the “brain,” acting as an annotator with a strong performance for contextualized output selection and correction. Our experiments show that LyricWhiz significantly reduces Word Error Rate compared to existing methods in English and can effectively transcribe lyrics across multiple languages. Furthermore, we use LyricWhiz to create the first publicly available, large-scale, multilingual lyrics transcription dataset with a CC-BY-NC-SA copyright license, based on MTG-Jamendo, and offer a human-annotated subset for noise level estimation and evaluation. We anticipate that our proposed method and dataset will advance the development of multilingual lyrics transcription, a challenging and emerging task. The code and dataset are available at https://github.com/zhuole1025/LyricWhiz.
@@ 1 Introduction
In this paper, we present LyricWhiz, a novel method for automatic lyrics transcription. LyricWhiz surpasses existing methods on various ALT datasets, resulting in a significant reduction in WER for English lyrics and providing accurate transcription results across multiple languages. Our system is robust, multilingual, and training-free. To achieve these results, we combined two powerful models from their respective domains as shown in Figure 1: Whisper, a weakly supervised speech transcription model, and GPT-4, a large language model (LLM) from the ChatGPT family. Whisper acts as the “ear” while GPT-4 serves as the “brain” by providing contextualized output selection and correction with strong performance [9]. We further use LyricWhiz to build a multilingual lyrics dataset, named MulJam, which is the first large-scale, multilingual lyrics transcription dataset without copyright-related issues.
@@ 2.3 Chat-based Large Language Models
c) The importance of a well-designed CoT [42], which effectively divides a complicated task into several phases and designs specific prompts for each phase, is widely acknowledged for enhancing LLM performance. We also proposes a concise CoT strategy, depicted in Section 3.2.
@@ 3.1 Whisper as Zero-shot Lyrics Transcriptor
Task: As a GPT-4 based lyrics transcription post-processor, your task is to analyze multiple ASR model-generated versions of a song’s lyrics and determine the most accurate version closest to the true lyrics. Also filter out invalid lyrics when all predictions are nonsense. Input: The input is in JSON format: {“prediction_1”: “line1;line2;…”, …} Output: Your output must be strictly in readable JSON format without any extra text: { “reasons”: “reason1;reason2;…”, “closest_prediction”: <key_of_prediction> “output”: “line1;line2…” } Requirements: For the "reasons" field, you have to provide a reason for the choice of the "closest_prediction" field. For the "closest_prediction" field, choose the prediction key that is closest to the true lyrics. Only when all predictions greatly differ from each other or are completely nonsense or meaningless, which means that none of the predictions is valid, fill in "None" in this field. For the "output" field, you need to output the final lyrics of closest_prediction. If the "closest_prediction" field is "None", you should also output "None" in this field. The language of the input lyrics is English.
@@ 3.3 Multilingual Lyrics Transcription Dataset
To prepare the dataset for training, it is essential to conduct line-level annotation. Timestamps can be obtained from the output of Whisper by aligning the lyrics both before and after ChatGPT processing. For the alignment of strings, the Levenshtein distance [43] is employed. To exclude aligned lines of lower confidence, the distance is normalized, setting a threshold at 0.2. The quality of annotation is further enhanced through two subsequent filtering stages. In the first stage, lines that exhibit unusually high character rates, exceeding 37.5 Hz, are eliminated. The second stage encompasses another Whisper iteration; segments yielding a transcription of "Thank you." are excluded. These segments, which typically represent instrumental sections, are believed to originate from Whisper’s training on data similar to video transcripts.
Following the construction process outlined above, we ultimately obtained a multilingual lyric transcription dataset, MulJam, consisting of 6,031 songs with 182,429 lines and a total duration of 381.9 hours. The dataset’s statistical information and comparisons with existing ALT datasets are presented in Table 2.
To our best knowledge, MulJam is the first publicly available large-scale dataset for multilingual lyrics transcription without copyright restrictions. While DALI [7] is another large-scale music dataset featuring multilingual lyrics, its restricted access and strict licensing requirements limit its applicability for downstream tasks. In contrast, MulJam is free from copyright-related constraints and can be utilized without approval, as the audio can be legally downloaded directly from public sources without the need for approval, making it easily accessible. This even includes audio that is permitted for use in the development of commercial software. Researchers are permitted to legally modify our dataset for derivative works and redistribution, provided they cite our work and adhere to the CC BY-NC-SA license. Furthermore, in contrast to the imbalanced language distribution in DALI, where English songs account for over $80\%$ of the total songs, our dataset includes a greater proportion of songs in other languages, which is advantageous for multilingual lyrics transcription.
@@ 5 Conclusion
6 Acknowledgements
We gratefully acknowledge the dataset post-processing work described in Section 3.3 offered by Jiawen Huang. Jiahao Pan and Wei Xue were supported by the Theme-based Research Scheme (T45-205/21-N) and Early Career Scheme (ECS-HKUST22201322), Research Grants Council of Hong Kong. Yinghao Ma is a research student at the UKRI Centre for Doctoral Training in Artificial Intelligence and Music, supported by UK Research and Innovation [grant number EP/S022694/1]. Yizhi Li is fully funded by an industrial PhD studentship (Grant number: 171362) from the University of Sheffield, UK.